Patriot Co. manufactures and sells three products: red, white, and blue. Their unit selling prices are red, $64; white, $94; and blue, $119. The per unit variable costs to manufacture and sell these products are red, $49; white, $69; and blue, $89. Their sales mix is reflected in a ratio of 5:4:2 (red:white:blue). Annual fixed costs shared by all three products are $159,000. One type of raw material has been used to manufacture all three products. The company has developed a new material of equal quality for less cost. The new material would reduce variable costs per unit as follows: red, by $6; white, by $16; and blue, by $6. However, the new material requires new equipment, which will increase annual fixed costs by $29,000.

Required:
1. Assume if the company continues to use the old material, determine its break-even point in both sales units and sales dollars of each individual product.
2. Assume if the company uses the new material, determine its new break-even point in both sales units and sales dollars of each individual product.
Answer #1
Red White Blue
1] Selling price per unit $              64.00 $               94.00 $             119.00
-Variable cost per unit $              49.00 $               69.00 $               89.00
Contribution margin per unit $              15.00 $               25.00 $               30.00
Sales mix 5 4 2
Weighted average unit contribution margin $              21.36
[(15*5+25*4+30*2)/(5+4+2)]
Break-even in sales units: Total
Break even in composite units = Annual fixed costs/Weighted average unit contribution margin = 159000/21.36 = 7444
Product wise units at BEP 3384 2707 1353 7444
[7444*5/11] [7444*4/11] [7444*2/11]
Breakeven in sales dollars:
Units at BEP*Selling price per unit = $        2,16,553 $         2,54,449 $         1,61,061 $    6,32,063
2] Red White Blue
Selling price per unit $              64.00 $               94.00 $             119.00
-Variable cost per unit $              43.00 $               53.00 $               83.00
Contribution margin per unit $              21.00 $               41.00 $               36.00
Sales mix 5 4 2
Weighted average unit contribution margin $              31.00
[(21*5+41*4+36*2)/(5+4+2)]
Break-even in sales units:
Break even in composite units = Annual fixed costs/Weighted average unit contribution margin = (159000+29000)/31 = 6065
Product wise units at BEP 2757 2205 1103 6065
[6065*5/11] [6065*4/11] [6065*2/11]
Breakeven in sales dollars:
Units at BEP*Selling price per unit = $        1,76,436 $         2,07,313 $         1,31,225 $    5,14,974